OpenStudy (anonymous):

A local highschool has 1250 students in grades 9-12 . Twenty-eight percent of the students in the school are in ninth grade. One-half of the ninth grade students ride the bus to school. how many ninth grade students ride the bus?

OpenStudy (texaschic101):

total 1250 students 28% are in 9th grade, (0.28)(1250) = 350 9th graders 1/2 ride the bus....1/2 of 350 = 350/2 = 175 175 9th graders ride the bus

OpenStudy (anonymous):

thank you so much.. I haven't done this type of problem in a while.. I completely forgot how,

OpenStudy (texaschic101):

glad to help :)

OpenStudy (anonymous):

I ought to just ask you from now on lol. (x nobody else will help me.. and i haven't done these problems in forever..

OpenStudy (texaschic101):

practice enough and it will all come back to you

OpenStudy (anonymous):

Always do c: again thank you (:

OpenStudy (texaschic101):

you got anymore that you need help with ?

OpenStudy (anonymous):

annd.. kinda yeah... x').. i have 5 left.. :c

OpenStudy (texaschic101):

ask away

OpenStudy (anonymous):

When Poll P asked students to identify their favorite kind of TV show , 12 students , 30% of the group, said their kind was sports. Poll Q asked 3 times as many as Poll P. how many students took Poll Q? (actual question: Amanda has this question on a test . Before she can answer it what information does she have to find first? A: The number of students who took the Poll P B: the number of students who took the Poll Q C: The number of students who picks sports in Poll Q D: The percent of Poll Q who chose sports.

OpenStudy (anonymous):

last year khat spent 1/4 ad 1/3 of her grooves income on her mortage payments.if kate spent $13470 on her mortage payments last year.what is grooves income last year?

OpenStudy (anonymous):

it was A.. c: next question :o?

OpenStudy (texaschic101):

poll P ... 12 students = 30% 12 = 30% of what 12 = .30x 12/.30 = x 40 = x 40 students in poll P 3 times as many in poll Q.....3(40) = 120 in poll Q oops...I did the whole problem....the answer is A, the number of people who took poll

OpenStudy (anonymous):

lol

OpenStudy (texaschic101):

ram.... 1/4 + 1/3 = 3/12 + 4/12 = 7/12 7/12 = 13470 7/12 of what is 13470 7/12x = 13470 x = 13470/(7/12) x = 13470 * 12/7 x = 161640/7 x = 23091.43 --> I think this is her income last year

OpenStudy (texaschic101):

anymore ?

OpenStudy (texaschic101):

oops...lost track of time...sorry but I got to go. I am sorry